AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 21 of 33  ELOG logo
ID Date Author Subject Text Attachments
  259   Wed Feb 23 19:22:55 2022 CN, TDcheck 

- Motor checked , OK

SW3
 10x 
  258   Wed Feb 23 18:58:12 2022 TD CNWednesday 23 February 15.00-23.00 shift

Assorted screenshots of the CARME, CRYRING
 6x 
  257   Wed Feb 23 16:57:35 2022 CB, CNOptimising motor motionOptimised motor motion during run R58

Moved 2 mm in (20.3 potentiometer)

29.47 potentiometer
  
  256   Wed Feb 23 10:56:06 2022 Laszlo, ThanassisCurrent status11:55
Temps OK
System Check OK
 10x 
  255   Wed Feb 23 09:50:19 2022 Laszlo, ThanassisMoving detector inWe are moving the detector in run 56   
  254   Wed Feb 23 08:58:30 2022 Laszlo, ThanassisCurrent status10:00
Temps OK
System Check OK
 10x 
  253   Wed Feb 23 07:58:38 2022 Laszlo, ThanassisCurrent status09:00
Temps OK
System Check OK
 10x 
  252   Wed Feb 23 06:58:19 2022 Laszlo, ThanassisCurrent status08:00
Temps OK
System Check OK
 11x 
  251   Wed Feb 23 06:00:25 2022 Jordan, GeorgeCurrent status07:00
Pressure Stable
Temps OK
 2022-02-23_07-06-36exRate.png2022-02-23_07-06-11exTotal.png2022-02-23_07-05-42xyTotal.png 
  250   Wed Feb 23 04:55:48 2022 Jordan, GeorgeCurrent status05:55
Temps OK
System Check OK
 2022-02-23_06-00-29xyTotal.png2022-02-23_05-59-36exRate.png2022-02-23_05-59-05exTotal.png 
  249   Wed Feb 23 03:52:13 2022 Jordan, GeorgeCurrent status04:52
Pressure seems to be climbing a little
but still stable (just peaked at end of check)
 2022-02-23_04-58-19-XYtotal.png2022-02-23_04-57-17-Ex-total-hist.png2022-02-23_04-56-36-Ex-rate-hist.pngpressures-1hour.jpgpressure-6hours.jpg 
  248   Wed Feb 23 02:52:09 2022 Jordan, GeorgeCurrent status03:52

Temps OK

System Checks OK
 2022-02-23_03-58-03xyTotal.png2022-02-23_03-57-34exTotal.png2022-02-23_03-56-52exRate.png 
  247   Wed Feb 23 01:56:59 2022 Jordan, GeorgeCurrent status02:56

Temps OK

write speed ~1.2Mb/s
 2022-02-23_03-01-54xyTotal.png2022-02-23_03-00-54exTotal.png2022-02-23_03-00-11exRate.png2022-02-23_03-10-45-aida1-1*L.png 
  246   Wed Feb 23 00:54:17 2022 Jordan, GeorgeCurrent status02:04

Temps OK

Data Rate - ~1.1Mb/s
 2022-02-23_02-02-53extotal.png2022-02-23_02-02-00lowEnergyXYtot.png2022-02-23_01-59-14exRate.png 
  245   Tue Feb 22 21:57:13 2022 George, JordanCurrent status23:53

Temperatures OK

System checks OK - white rabbit
 2022-02-22_23-53-38-lowenergy-xytotal.png2022-02-22_23-50-45-Ex-rate-hist.png2022-02-22_23-49-20-Ex-total-hist.png20220223_002946.jpg 
  244   Tue Feb 22 21:09:39 2022 David, YuriCurrent statusPressure seems as before 19:10.

Merger, Tape server (WS3)
checks: DONE, writing ~1 MB/s, refreshing
 2022-02-22_22-12-55xytotal.png2022-02-22_22-12-21exrate.png2022-02-22_22-12-00extotal.png 
  243   Tue Feb 22 20:08:40 2022 David, YuriCurrent statusRecently the pressure started to raise
a bit, we have increased the temperature
from 129.1 K to 129.2 K. Current pressure
 2022-02-22_21-14-23exrate.png2022-02-22_21-14-01xytotal.png2022-02-22_21-13-11extotal.png 
  242   Tue Feb 22 19:28:35 2022 David, YurivacuumAt 19:10 Frank called us that vacuum in
sections 08 and 10 got worse by about 1.5
times. There is a possible danger that vacuum
  
  241   Tue Feb 22 19:11:18 2022 DavidCurrent statusStill trying to recover previous pressures,
more details in the next entry.

Merger, Tape server (WS3)
 2022-02-22_20-16-26eyrate.png2022-02-22_20-15-34.eytotal.png 
  240   Tue Feb 22 18:23:02 2022 David. YuryVacuum problemAt 19:10 the vacuum has increased suddenly
everywhere in the ring, roughly a factor
1.5.
  
ELOG V3.1.4-unknown